OpenStudy (anonymous):

solve this, 16 raise to 4-3x= (.25) raise to x+2

OpenStudy (anonymous):

16 is 4^(2) and 0.25 is 4^(-1) so 4^(2(4-3x)) = 4^(-1(x+2)) so 8-6x = -x-2 so 10 = 5x therefore x = 2

OpenStudy (anonymous):

i agree with your ans!!!!!!!